OpenStudy (anonymous):

cos²β(1+tan²β)=1. Show this is an identity. I have so much trouble with these....

OpenStudy (anonymous):

It follows from \[1+\tan^2\beta = \sec^2\beta\] (which in turn follows from dividing the identity \[\sin^2\beta + \cos^2\beta = 1 \text{ by } \cos^2\beta \]

OpenStudy (anonymous):

I still don't see it. what would i do next?

OpenStudy (anonymous):

Wait, it's easier. Just multiply out the equation you have to start and it follows immediately.

OpenStudy (anonymous):

tan^2 = sin^2/cos^2 => tan^2 x cos^2 = sin^2

OpenStudy (anonymous):

I'm still confused!

OpenStudy (anonymous):

Oh dear :(